OpenStudy (anonymous):

[6.03] What is the value of the y variable in the solution to the following system of equations? 5x + 3y = 7 3x - 5y = -23

OpenStudy (anonymous):

do u know what the x-value is

OpenStudy (anonymous):

no i dont and i really need help i have an half hour to turn this in

OpenStudy (anonymous):

y=1/3(7-5x)

OpenStudy (anonymous):

multiply the tope equation by(3/5) then subtract bottom from top to cancel out x then use that to solve for y and then plug y value back into one of the equations to solve for x know what I mean?

OpenStudy (anonymous):

top**

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so you would get 3x + (9/5)y = (21/5) -3x - 5y = -23 (34/5)y = (136/5) y = 136/34 y = 4 3x - 5(4) = -23 x = -3/3 x = -1

OpenStudy (anonymous):

Do you see what I did?